OpenStudy (anonymous):

the radius is 9 centimeters. which is the closest area of the circles a.54 cm 2 b.245cm2 c.27 cm 2 d.18 cm 2

OpenStudy (mathmale):

The area of a circle is \[\pi r^2, so what do you do with the given radius? What happens when you multiply the result by pi, whether you use 22/7 or 3.14 or 3.1459?

OpenStudy (anonymous):

????????????????????? ._.

OpenStudy (mathmale):

The area of a circle is A = pi*r^2 (pi times the square of the radius). Therefore, when the radius is 9 cm, the square of the radius is 81 cm^2. Multiplying this by 22/7 (which is an approximation for pi), \[\frac{ 22*81 cm^2 }{ 7 },\] which is approximately what value? And this value is closest to which of the four possible answers given?

OpenStudy (anonymous):

area=pir^2 area=3.14*81 area= ??